# Numerical study on wave attenuation via 2D fully kinetic electromagnetic particle-in-cell simulations

### v1

@@ Abstract
The propagation and absorption of electromagnetic waves in plasma is one of the fundamental issues in plasma physics. The electromagnetic particle-in-cell method with the finite-difference time-domain solver plus Monte Carlo collision model would be the most accurate method to simulate the wave-plasma interaction. However, the numerical effects of this method have not been carefully investigated especially in two dimensions. In this paper, the 2D PIC method is used to study the electromagnetic wave attenuation by fluorescent lamp plasma tubes. The study finds that the number of macro-particles and the incident electromagnetic wave amplitude have minor effects on the wave attenuation within a certain appropriate parameter range. Furthermore, the effects of electromagnetic wave frequency, the plasma distribution structures, and collision types on wave attenuation are investigated. Particularly, it is found that the staggered way of arranging the plasma tubes can achieve better wave attenuation than the parallel way, which agrees with our recent experimental observation.
@@ 1 Introduction
Plasma, a state of matter characterized by relatively high temperatures and energy, consists of charged ions and electrons. It plays a crucial role in the interaction with electromagnetic waves in various scientific and technological fields. As electromagnetic waves propagate through plasma, they experience refraction, reflection, and attenuation. These characteristics are useful in the development of materials for absorbing electromagnetic waves, antenna design, and wireless communication modulation techniques[1, 2]. The study of plasma and electromagnetic waves has revealed its physical properties, offering broad application prospects in fields such as telecommunications.
@@
Different computational methods are suitable for different object orientations. As a representative analytical method, the WKB (Wentzel Kramers Brillouin) method is a well-established approximate calculation technique for solving electromagnetic wave propagation problems in slowly varying media, and it is also one of the most commonly used methods when discussing the interaction between plasmas and electromagnetic waves. In 2003, Shaobin Liu and his colleagues employed the WKB method to study the influence of electron density, collision frequency, and electromagnetic wave frequency on the attenuation of electromagnetic waves in plasmas with linear and exponential distributions[3, 4]. The WKB method, which treats stratified plasmas through geometric optics approximation and requires each layer to be uniform with minimal parameter variations, is applicable to solving electromagnetic wave propagation in one-dimensional plasmas limited in simple structures.
@@
When the simulation object is two-dimensional or even three-dimensional, the main applicable methods are the finite-difference time-domain (FDTD) method. The principle of the finite-difference time-domain (FDTD) method involves discretizing the electric field strength $E$ components and magnetic field strength H components in space and time using a staggered grid. Each $E$ or $H$ component is surrounded by four $H$ or $E$ components, respectively. By analyzing the surrounding components and using this discretization approach, the Maxwell curl equations with time variables are transformed into a set of difference equations, which are then solved step by step on the time axis to obtain the spatial electromagnetic fields. Throughout the development of FDTD, various studies have been conducted on the FDTD method for different applications [5, 6, 7]. However, the FDTD method uses complex permittivity to approximate all plasma behaviors, which cannot resolve detailed particle motion and trajectories. In contrast, the particle-in-cell (PIC) method combines the FDTD method to solve the electromagnetic fields and tracks the motion of plasma particles by solving the equations of motion. In the PIC method, the plasma is naturally represented in the form of macro-particles and is connected to the Maxwell’s equations through the charge and current density deposited by the particles at grid points. Moreover, collisions between charged particles can be handled using the Monte Carlo collision (MCC) model. Therefore, if computational resources permit, using PIC combined with the MCC model and the FDTD solver would be the most accurate method for studying the interaction between electromagnetic waves and plasmas.
@@
From existing theoretical and experimental research, it is known that to achieve the best absorption effect of electromagnetic waves, it is necessary to allow the waves to enter the plasma as much as possible and maximize the propagation distance of the waves within the plasma to achieve the lowest reflection of the electromagnetic signal. To enhance the plasma’s absorption of electromagnetic waves, scholars have conducted studies on the transmission performance of electromagnetic waves through plasma with different spatial arrangements. Vidmar[12] discovered in 1990 that atmospheric pressure air or helium plasma can absorb electromagnetic waves across a wide frequency band (from very high frequency to X-band), and indicated that the absorption effect is better when the plasma density is distributed inversely proportional to the distance from the plasma source. This research provides a certain reference for the plasma density distribution in array arrangement. Xuesong Deng from National University of Defense Technology [13] achieved different electromagnetic wave absorption effects by combining a plasma array with absorbent materials and adjusting the electron density within each discharge tube to change the distribution state of the plasma array. The Cappelli team at Stanford University arranged gas discharge tubes in a periodic three-dimensional “stake” structure to form a tunable plasma photonic crystal. Their research concluded that the transmission effect with all discharge tubes open is different from the superposition of the transmission effects in the vertical and horizontal directions, demonstrating that the influence of different spatial structural positions on electromagnetic wave transmission is not the same and is mutually coupled[14].
@@
In the aforementioned work applying the Particle-in-Cell (PIC) method to study wave-plasma interactions, the PIC method is directly used as a tool for investigating physical problems. However, the detailed numerical impact of PIC parameters on wave-plasma interactions has not been reported or deeply explored. Simultaneously, studies on plasma arrays with simple configurations have demonstrated that the arrangement structure of the plasma has a significant effect on the attenuation efficiency, capable of enhancing absorption. A previous work of us in 2024, Yan et al. [15], studied the effects of numerical and physical parameters, such as the electromagnetic wave frequency, amplitude, plasma temperature, thickness, and collision types on wave attenuation through one dimensional fully dynamic electromagnetic PIC simulations. However, only the absorption effects of one-dimensional distributed plasma have been investigated, lacking a two-dimensional PIC study on the absorption of electromagnetic waves by different plasma distributions. Therefore, in this paper, we employ a two-dimensional electromagnetic PIC with Monte Carlo Collisions (MCC) to study the wave-plasma interaction problem. In Sec.2, the numerical methods are described, including the applied PIC code and considered simulation setup, the characteristic simulation results and diagnostics , and the numerical effects of PIC parameters are examined. In Sec.3, the influence of numerical and physical parameters is studied, including the number of macro-particles, EM wave frequency, amplitude, collisions, and plasma distribution structures. At last in Sec.4, conclusions are drawn and future research plans are described.
@@ 2.1 The applied PIC code and algorithms
In this study, the simulation tool we employed is the WarpX open-source PIC code [16], a highly parallelized and optimized code capable of scaling on some of the world’s largest supercomputers, and in $2022$, it was awarded the ACM Gordon Bell Prize. Although WarpX was initially developed for plasma-based accelerator applications [17, 18], it has now been extended to a variety of plasma simulation fields, including strong-field quantum electrodynamics [19], relativistic magnetic reconnection [20], plasma propulsion [21, 22], and more. WarpX features a range of capabilities, supporting $1-3$D Cartesian and $2$D cylindrical (RZ) coordinate systems, and includes electromagnetic field solvers such as FDTD (Finite-Difference Time-Domain), CKC (Current-Conserving, Charge-Conserving algorithm), PSATD (Pseudo-Spectral Analytical Time-Domain), and MLMG (Multi-Level Multi-Grid) electrostatic field solver. Additionally, it integrates modules for Coulomb collisions and ionization Monte Carlo collisions, supports adaptive mesh refinement, Perfectly Matched Layers (PML) and other absorbing boundary conditions, Lorentz boosted frame techniques, embedded boundary handling, and efficient CPU/GPU parallel domain decomposition techniques. To study the process of interaction between incident electromagnetic waves and plasma, as well as the interaction of these waves with electrons and ions in the plasma, we selected the $2$D geometric configuration of WarpX, employed the classical FDTD electromagnetic field solver, and Perfectly Matched Layers absorbing boundary and the Monte Carlo ionization collision algorithm. These settings enable us to self-consistently resolve the interaction between the incident electromagnetic waves and those induced by the plasma, as well as their interactions with electrons and ions of the plasma.
Figure 1: Illustration of the simulation setup.
@@ 2.2 Simulation setup
The fluorescent lamp plasma tube, as a commonly used plasma generation device, is readily accessible and offers relatively simple control of plasma parameters. After years of development, the fluorescent lamp plasma tube has garnered extensive research attention and has achieved significant accomplishments in the field [9, 23, 24]. Many companies and research institutions continuously improve these devices for economic purposes, aiming to ionize the maximum number of neutral atoms with the lowest power, making it the optimal choice for the engineering of discharge devices. Considering the range of plasma parameters, the degree of reliability, and the controllability of the discharge state, as well as the ease of constructing an array of devices, and referring to the existing literature on such devices, this paper selects the plasma discharge tube in a closed discharge system as the two-dimensional simulation geometry.
@@
In the plasma region between $z_{1}$ and $z_{2}$, the particle number density is $n_{0}=n_{e}=n_{i}$, with particles distributed uniformly and randomly. The velocity distributions of electrons and ions follow the Maxwell-Boltzmann distribution, with temperatures of $T_{e}$ and $T_{i}$, respectively. The subscripts ‘$e$’ and ‘$i$’ denote electrons and ions, while ‘$a$’ represents neutral atoms. The number density of neutral atoms is $n_{a}$, and their temperature is $T_{a}$. These parameters are used in the Monte Carlo Collision (MCC) algorithm to handle the collision processes between particles. In this study, no additional boundaries are required to be set for the plasma region, as we found that the ion motion has a
Figure 2: The propagation process of electromagnetic waves.
negligible effect on the simulation results when the electromagnetic wave passes through the plasma, while the electrons are effectively confined within the plasma region. The ion species considered is argon, and the MCC algorithm includes the processes electron-neutral scattering, excitation and ionization for argon, which were conformed to be the most important collision types in our previous 1D study[15].
2.3 Characteristic simulation results and diagnoses
[…]
(a) The transmitted power during the first half-wave period.
(b) The transmitted power during the second half-wave period.
Figure 3: The Poynting Vector at $z_{t}=800$.
@@ 3.1 The impact of macro-particles number
Firstly, we examine the impact of the number of macro-particles on the simulation. We investigate the effect of the total number of macro-particles $n_{pp}$ on the attenuation of electromagnetic waves in the simulation by varying $n_{pp}$ to $10^{6}$, $10^{7}$, and $10^{8}$, while keeping all other conditions constant. As shown in the left panel of Fig.5, the incident power does not change with the variation in the number of macro-particles, and the lines for $n_{pp}=10^{7}$ and $n_{pp}=10^{8}$ overlap when reflecting, with only a slight difference observed for the data with $n_{pp}=10^{6}$. The transmitted power through the plasma for the three different particle numbers is shown in the right panel of Fig.5. The lines for $n_{pp}=10^{7}$ and $n_{pp}=10^{8}$ are still highly overlapping, while a difference at the peak is observed for $n_{pp}=10^{6}$ compared to the other two particle numbers.
@@ 3.2 Varying EM wave frequency
Figure 7: wave attenuation over EM wave frequency.
As shown in the Fig.7, for a plasma density of $10^{18}$m-3and a plasma frequency of $9$ GHz, the relationship between electromagnetic wave attenuation and the frequency of the wave is depicted for different incident electromagnetic wave frequencies. We can observe that when the frequency of the incident electromagnetic wave is low, the attenuation is minimal, particularly the $\textrm{Att}_{\textrm{S}}^{\textrm{R}}$line after subtracting the reflected wave is more pronounced. This is because when the electromagnetic wave frequency is less than the plasma frequency, causing total reflection of the wave at the plasma interface, preventing it from entering the plasma interior. As the electromagnetic wave frequency approaches $9$ GHz, reflection decreases, and the amount of electromagnetic wave attenuation significantly increases. This is due to the plasma frequency being close to the frequency of the incident wave at this point, causing the wave to propagate near the cutoff region and leading to resonant absorption, which can cause substantial attenuation of the electromagnetic wave. Continuing to increase the frequency of the electromagnetic wave moves it further from the cutoff frequency, and it can be seen that the $\textrm{Att}_{\textrm{S}}^{\textrm{R}}$ and $\textrm{Att}_{\textrm{S}}$ lines gradually converge, with reflection gradually diminishing. The attenuation of the electromagnetic wave by the plasma will then primarily be due to collisional absorption. When the electromagnetic wave frequency is too high, the electrons in the plasma cannot respond quickly enough to the changes in the wave, and they cannot polarize effectively. As a result, their ability to absorb the energy of the electromagnetic wave is greatly reduced, leading to decreased attenuation of the wave within the plasma, making it easier for the wave to penetrate through the plasma.
3.3 Varying EM wave amplitude
Figure 9: Wave attenuation over EM wave amplitdue.
3.4 Plasma distribution structures
This section discusses the impact of different plasma distribution structures on electromagnetic wave attenuation. As shown in the Fig.11, the simulation images depict various distribution states with an amplitude of $2×10^{5}$/m,
@@ 3.4 Plasma distribution structures
an electromagnetic wave frequency of $15$ GHz, and a certain radius. Fig.11a shows a staggered two-row distribution where the line connecting the centers of adjacent three discharge tubes can form an equilateral triangle, Fig.11b shows a two-row parallel plasma distribution, and Fig.11c shows a single-row plasma distribution. Additionally, in order to reduce the computational load and simulate the absorption effect of a large number of discharge tubes on electromagnetic waves, periodic boundary conditions are applied to both the upper and lower boundaries of the calculation area. This ensures that the plasma region continues to exist outside the calculation area, thereby eliminating the phenomenon of electromagnetic waves bypassing at the farthest boundary due to an excessively large wavelength.

### v2

@@ Abstract
The propagation and absorption of electromagnetic waves in plasma is one of the fundamental issues in plasma physics. The electromagnetic particle-in-cell method with the finite-difference time-domain solver plus Monte Carlo collision model would be the most accurate method to simulate the wave-plasma interaction. However, the numerical effects of this method have not been carefully investigated especially in two dimensions. In this paper, the 2D PIC method is used to study the electromagnetic wave attenuation by fluorescent lamp plasma tubes. The study finds that the number of macro-particles and the incident electromagnetic wave amplitude have minor effects on the wave attenuation within a certain appropriate parameter range. Furthermore, the effects of electromagnetic wave frequency, the plasma distribution structures, and collision types on wave attenuation are investigated. Particularly, it is found that the staggered way of arranging the plasma distribution structures can achieve better wave attenuation than the parallel way, which agrees with our recent experimental observation.
@@ 1 Introduction
Plasma plays a crucial role in the interaction with electromagnetic waves in various scientific and technological fields. As electromagnetic waves propagate through plasma, they experience refraction, reflection, and attenuation. These characteristics are useful in the development of materials for absorbing electromagnetic waves, antenna design, and wireless communication modulation techniques[1, 2]. The study of plasma and electromagnetic waves has revealed its physical properties, offering broad application prospects in fields such as telecommunications.
@@
In the $1950$s and $1960$s, driven by military needs, both domestically and internationally, a substantial amount of theoretical research was conducted on the interaction between plasma and electromagnetic waves. Over the decades, significant progress has been made in theoretical research. Ginzburg, Vidmar, and others established a series of theoretical models for the interaction between electromagnetic waves and plasma [3, 4], which have been widely applied by subsequent scholars. The development of computers has also led to great advancements in the numerical simulation of the interaction between plasma and electromagnetic waves. In $2003$, Shaobin Liu and his colleagues employed the WKB method to study the influence of electron density, collision frequency, and electromagnetic wave frequency on the attenuation of electromagnetic waves in plasmas with linear and exponential distributions[5, 6]. The WKB method, which treats stratified plasmas through geometric optics approximation and requires each layer to be uniform with minimal parameter variations, is applicable to solving electromagnetic wave propagation in one-dimensional plasmas limited in simple structures.
@@
When the simulation object is two-dimensional or even three-dimensional, the main applicable methods are the finite-difference time-domain (FDTD) method. In $2009$, Chaudhury and his team used the Finite-Difference Time-Domain (FDTD) method to study the Radar Cross Section (RCS) of a metal plate covered with a plasma with three-dimensional inhomogeneity. The frequency of the incident electromagnetic wave was in the range of $3$-$9$ GHz. They found that the attenuation of the electromagnetic wave by the plasma was related to the angle of incidence[7, 8, 9]. However, the FDTD method uses complex permittivity to approximate all plasma behaviors, which cannot resolve detailed particle motion and trajectories. In contrast, the particle-in-cell (PIC) method combines the FDTD method to solve the electromagnetic fields and tracks the motion of plasma particles by solving the equations of motion. In the PIC method, the plasma is naturally represented in the form of macro-particles and is connected to the Maxwell’s equations through the charge and current density deposited by the particles at grid points. Moreover, collisions between charged particles can be handled using the Monte Carlo collision (MCC) model. Therefore, if computational resources permit, using PIC combined with the MCC model and the FDTD solver would be the most accurate method for studying the interaction between electromagnetic waves and plasmas.
Figure 1: Illustration of the simulation setup.
@@
From existing theoretical and experimental research, we know that to achieve the optimal absorption effect of electromagnetic waves, it is necessary to allow electromagnetic waves to enter the plasma as much as possible and maximize the propagation distance of electromagnetic waves within the plasma to achieve minimal reflection of electromagnetic signals. In order to enhance the high absorption of plasma for electromagnetic waves, as well as to meet the demands of increasingly advanced radar detection technology and communication control technology, scholars have studied the transmission performance of electromagnetic waves in plasmas with different spatial arrangements[3, 14].
Figure 2: The propagation process of electromagnetic waves.
The fluorescent lamp plasma tube, as a commonly used plasma generation device, is readily accessible and offers relatively simple control over plasma parameters. After years of development, the fluorescent lamp plasma tube has received extensive attention in the field of wave attenuation research and has achieved significant accomplishments[11, 15, 16]. For example, Danilov et al[17]. studied the reflection, transmission, and absorption coefficients of periodically arranged plasmas for electromagnetic waves in the $4$-$12$ GHz range. Howlader et al[18]. investigated the attenuation of $10$ GHz electromagnetic waves by fluorescent tube plasma and used the microwave interferometry method to determine that the electron number density of the fluorescent lamp plasma ranges from $3.5\times 10^{10}\text{ to }2.6\times 10^{11}\text{ cm}^{-3}$, with collision frequencies between $7$ and $40$ GHz. The Cappelli team at Stanford University arranged gas discharge tubes in a periodic three-dimensional “stake” structure to form a tunable plasma photonic crystal. Their research concluded that the transmission effect with all discharge tubes open is different from the superposition of the transmission effects in the vertical and horizontal directions, demonstrating that the influence of different spatial structural positions on electromagnetic wave transmission is not the same and is mutually coupled[19]. Although some progress has been made in the study of wave attenuation in plasma discharge tube arrays, to our knowledge, there are few papers that simulate and study the attenuation of electromagnetic waves in plasma discharge tubes with different structural distributions using 2D PIC methods. Therefore, considering the range of plasma parameters, reliability, controllability of the discharge state, cost-effectiveness, and the convenience of constructing equipment arrays, and referring to the existing literature on such devices, this paper selects the plasma discharge tube as the two-dimensional simulation geometry.
@@
In the aforementioned work applying the Particle-in-Cell (PIC) method to study wave-plasma interactions, the PIC method is directly used as a tool for investigating physical problems. However, the detailed numerical impact of PIC parameters on wave-plasma interactions has not been reported or deeply explored. Simultaneously, studies on plasma arrays with simple configurations have demonstrated that the arrangement structure of the plasma has a significant effect on the attenuation efficiency, capable of enhancing absorption. A previous work of us in $2024$, Yan et al. [20], studied the effects of numerical and physical parameters, such as the electromagnetic wave frequency, amplitude, plasma temperature, thickness, and collision types on wave attenuation through one dimensional fully dynamic electromagnetic PIC simulations. However, only the absorption effects of one-dimensional distributed plasma have been investigated, lacking a two-dimensional PIC study on the absorption of electromagnetic waves by different plasma distributions.
As far as we know, there are currently few research papers on the study of electromagnetic wave attenuation in plasma using two-dimensional ($2$D) Particle-in-Cell (PIC) simulations. Therefore, in this paper, we employ a two-dimensional electromagnetic PIC with Monte Carlo Collisions (MCC) to study the wave-plasma interaction problem. In Sec.2, the numerical methods are described, including the applied PIC code and considered simulation setup, the characteristic simulation results and diagnostics , and the numerical effects of PIC parameters are examined. In Sec.3, the influence of numerical and physical parameters is studied, including the number of macro-particles, EM wave frequency, amplitude, collisions, and plasma distribution structures. At last in Sec.4, conclusions are drawn and future research plans are described.
@@ 2.1 The applied PIC code and algorithms
In this study, the simulation tool we employed is the WarpX open-source PIC code [21], a highly parallelized and optimized code capable of scaling on some of the world’s largest supercomputers, and in $2022$, it was awarded the ACM Gordon Bell Prize. Although WarpX was initially developed for plasma-based accelerator applications [22, 23], it has now been extended to a variety of plasma simulation fields, including strong-field quantum electrodynamics [24], relativistic magnetic reconnection [25], plasma propulsion [26, 27], and more. WarpX features a range of capabilities, supporting $1-3$D Cartesian and $2$D cylindrical (RZ) coordinate systems, and includes electromagnetic field solvers such as FDTD (Finite-Difference Time-Domain), CKC (Current-Conserving, Charge-Conserving algorithm), PSATD (Pseudo-Spectral Analytical Time-Domain), and MLMG (Multi-Level Multi-Grid) electrostatic field solver. Additionally, it integrates modules for Coulomb collisions and ionization Monte Carlo collisions, supports adaptive mesh refinement, Perfectly Matched Layers (PML) and other absorbing boundary conditions, Lorentz boosted frame techniques, embedded boundary handling, and efficient CPU/GPU parallel domain decomposition techniques.
(a) The transmitted power during the first half-wave period.
(b) The transmitted power during the second half-wave period.
Figure 3: The Poynting Vector at $z_{t}=800$.
To study the process of interaction between incident electromagnetic waves and plasma, as well as the interaction of these waves with electrons and ions in the plasma, we selected the $2$D geometric configuration of WarpX, employed the classical FDTD electromagnetic field solver, and Perfectly Matched Layers absorbing boundary and the Monte Carlo ionization collision algorithm. These settings enable us to self-consistently resolve the interaction between the incident electromagnetic waves and those induced by the plasma, as well as their interactions with electrons and ions of the plasma.
@@ 2.2 Simulation setup
In the plasma region between $z_{1}$ and $z_{2}$, the particle number density is $n_{0}=n_{e}=n_{i}$, with particles distributed uniformly and randomly. The velocity distributions of electrons and ions follow the Maxwell-Boltzmann distribution, with temperatures of $T_{e}$ and $T_{i}$, respectively. The subscripts ‘$e$’ and ‘$i$’ denote electrons and ions, while ‘$a$’ represents neutral atoms. The number density of neutral atoms is $n_{a}$, and their temperature is $T_{a}$. These parameters are used in the Monte Carlo Collision (MCC) algorithm to handle the collision processes between particles. In this study, no additional boundaries are required to be set for the plasma region, as we found that the ion motion has a negligible effect on the simulation results when the electromagnetic wave passes through the plasma, while the electrons are effectively confined within the plasma region. The ion species considered is argon, and the MCC algorithm includes the processes electron-neutral scattering, excitation and ionization for argon, which were conformed to be the most important collision types in our previous 1D study[20].
2.3 Characteristic simulation results and diagnoses
[…]
The simulation was run for $20000$ time steps. Most parameters were chosen based on reference[20, 28]. $E_{0}=5\times 10^{4}$ V/m was selected such that $E_{0}/n_{a}\approx 208$ Td, which is within a reasonable range[29]. The simulation was conducted on a personal desktop computer equipped with an Intel® Xeon® Platinum 8269CY CPU @ $2.50$GHz. This computer possesses $26$ physical cores and $52$ logical processors. Throughout the simulation, $16$ MPI (Message Passing Interface) processes were employed to parallelize the simulation, aiming to enhance the efficiency and speed of the simulation.
Figure 7: wave attenuation over EM wave frequency.
@@ 3.1 The impact of macro-particles number
Firstly, we examine the impact of the number of macro-particles on the simulation. We investigate the effect of the total number of macro-particles $n_{pp}$ on the attenuation of electromagnetic waves in the simulation by varying $n_{pp}$ to $10^{6}$, $10^{7}$, and $10^{8}$, while keeping all other conditions constant. As shown in the left panel of Fig.5, the incident power does not change with the variation in the
Figure 9: Wave attenuation over EM wave amplitdue.
number of macro-particles, and the lines for $n_{pp}=10^{7}$ and $n_{pp}=10^{8}$ overlap when reflecting, with only a slight difference observed for the data with $n_{pp}=10^{6}$. The transmitted power through the plasma for the three different particle numbers is shown in the right panel of Fig.5.The lines for $n_{pp}=10^{7}$ and $n_{pp}=10^{8}$ are still highly overlapping, while a difference at the peak is observed for $n_{pp}=10^{6}$ compared to the other two particle numbers.
@@ 3.2 Varying EM wave frequency
As shown in the Fig.7, for a plasma density of $10^{18}$m-3and a plasma frequency of $9$ GHz, the relationship between electromagnetic wave attenuation and the frequency of the wave is depicted for different incident electromagnetic wave frequencies. We can observe that when the frequency of the incident electromagnetic wave is low, the attenuation is minimal, particularly the $\textrm{Att}_{\textrm{S}}^{\textrm{R}}$line after subtracting the reflected wave is more pronounced. This is because when the electromagnetic wave frequency is less than the plasma frequency, causing total reflection of the wave at the plasma interface, preventing it from entering the plasma interior. As the electromagnetic wave frequency approaches $9$ GHz, reflection decreases, and the amount of electromagnetic wave attenuation significantly increases.
@@
This is due to the plasma frequency being close to the frequency of the incident wave at this point, causing the wave to propagate near the cutoff region and leading to resonant absorption, which can cause substantial attenuation of the electromagnetic wave. Continuing to increase the frequency of the electromagnetic wave moves it further from the cutoff frequency, and it can be seen that the $\textrm{Att}_{\textrm{S}}^{\textrm{R}}$ and $\textrm{Att}_{\textrm{S}}$ lines gradually converge, with reflection gradually diminishing. The attenuation of the electromagnetic wave by the plasma will then primarily be due to collisional absorption. When the electromagnetic wave frequency is too high, the electrons in the plasma cannot respond quickly enough to the changes in the wave, and they cannot polarize effectively. As a result, their ability to absorb the energy of the electromagnetic wave is greatly reduced, leading to decreased attenuation of the wave within the plasma, making it easier for the wave to penetrate through the plasma.
3.3 Varying EM wave amplitude
3.4 Plasma distribution structures
This section discusses the impact of different plasma distribution structures on electromagnetic wave attenuation. As shown in the Fig.11, the simulation images depict various distribution states with an amplitude of $2$×$10^{5}$/m, an electromagnetic wave frequency of $15$ GHz, and a certain radius. Fig.11a shows a staggered two-row distribution where the line connecting the centers of adjacent three discharge tubes can form an equilateral triangle, Fig.11b shows a two-row parallel plasma distribution, and Fig.11c shows a single-row plasma distribution. Additionally, in order to reduce the computational load and simulate the absorption effect of a large number of discharge tubes on electromagnetic waves, periodic boundary conditions are applied to both the upper and lower boundaries of the calculation area. This ensures that the plasma region continues to exist outside the calculation area, thereby eliminating the phenomenon of electromagnetic waves bypassing at the farthest boundary due to an excessively large wavelength.